OpenStudy (anonymous):

Completely factor the difference of cubes 1-x^27

OpenStudy (anonymous):

\[a^2-b^3=(a-b)(x^2+ab+b^2)\] in your case \(a=1,b=x\)

OpenStudy (anonymous):

no, in your case \(a=1,b=x^3\)

OpenStudy (loser66):

* b = x^9

OpenStudy (anonymous):

yeah one of those...

OpenStudy (anonymous):

so is it suppose to look like this? 1-x^27=(1-x^9)(1+x+x^3)

OpenStudy (anonymous):

could anyone help me with my question

OpenStudy (anonymous):

During his 2009 MVP season for the St. Louis Cardinals, Albert Pujols hit 93 singles, 45 doubles, 1 triple, and 47 home runs in a total of 568 at bats. The following table arranges these data in terms of the number of bases for a hit, counting 0 bases for the times when he did not get a hit and 4 bases for home runs. Number of Bases | Number of Times 0 | 382 1 | 93 2 | 45 3 | 1 4 | 47 ............ it wants to know his typiical bat in 2009 to the nearest thousandth

OpenStudy (agent0smith):

I hope for your sake you don't have to completely factor it http://www.wolframalpha.com/input/?i=completely+factor+1-x%5E27 maybe you just have to do it as a difference of two cubes and be done.